OpenStudy (anonymous):

Electrical energy (E) is measured in kilowatt hours (kwh) using the formula E = P * t , where P is the power in kilowatts and t is the time in hours. If 120 kwh of electrical energy are created over a period of 15 hours, how many kilowatts of power are used each hour? A-135 B-1,800 C-8 D-105

OpenStudy (anonymous):

same thing as the other one you just substitute and solve for E i believe

OpenStudy (texaschic101):

its a matter of substitution... E = P * T E = 120 * 15 ...you can do the math

OpenStudy (anonymous):

I got B.

OpenStudy (anonymous):

yes that is correct

OpenStudy (anonymous):

sorry today im having a total brain fart. thank you so much

OpenStudy (anonymous):

your welcome that happens to me all the time

OpenStudy (anonymous):

that is wrong it is not 1800... what did i do wrong? @jewelshipp

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

it was 8. how is that even possible